### v1

# ISAC channel simulator GBSM

Abstract
Integrated Sensing and Communication (ISAC) is one of the key technologies in 6G, and related research and standardization efforts are progressing vigorously. Wireless channel simulation is the cornerstone for the evaluation and optimization of wireless communication technologies. This paper proposes a design and implementation method for an ISAC channel simulation based on a Geometry-Based Stochastic Model (GBSM) simulation framework. First, we introduce the progress of 3GPP ISAC channel standardization and the key topics of discussion. Second, addressing the current lack of a standardized ISAC channel simulation framework, we propose a cascaded ISAC channel simulation framework based on GBSM, leveraging our team’s related measurements, analyses, and proposal results. Based on this framework, we develop and design the ISAC channel simulator BUPTCMCC-6G-CMG+. Finally, we analyze and validate the simulation platform results, and provide some prospects for future ISAC testing research combined with channel simulators.
keywords:
ISAC standardization, channel simulator, GBSM, concatenate, ISAC testing.
1 Introduction
The next-generation mobile communication technology (6G) is expected to advance traditional communication capabilities such as high capacity, low latency, and high reliability while also achieving native intelligence, integrated sensing and communication (ISAC), and ubiquitous connectivity. The boundary between the physical and digital worlds will become increasingly blurred, providing users with a ubiquitous and immersive experience[1, 2]. ISAC technology is anticipated to drive the realization of various use cases, especially in target detection, localization, tracking, imaging, and estimation of target type, speed, and angle. These capabilities will serve applications in autonomous driving, industrial automation, digital twin reconstruction of physical environments, gesture detection, and activity recognition[3, 4, 5]. The sensing information obtained by ISAC can enable the construction of smart homes, smart factories, smart cities, and intelligent transportation systems. Additionally, ISAC technology introduces new performance dimensions compared to traditional communication technologies, such as detection probability, speed, angle, delay, sensing accuracy, and resolution, as well as sensing capacity density[6].
The transmission medium at both the transmitter and receiver determines the performance limits of wireless communication systems. The wireless channel is a critical component in the research and evaluation of wireless communication technologies[7].The evaluation scenarios for ISAC present new key requirements for channel models in terms of sensing modes, target modeling, and modeling methods. Firstly, in terms of sensing modes, traditional channel models are bi-static channels for communication. Sensing scenarios require consideration of mono-static channel models for self-sensing, thus ISAC channel models need to consider both mono-static and bi-static sensing channel models. Secondly, in target modeling for sensing, ISAC applications add the evaluation of target detection, localization, identification, and tracking, such as intrusion detection, UAV and V2X tracking, and recognition of different types and sizes of targets[8]. This necessitates leveraging the scattering properties, polarization characteristics, and Doppler information of electromagnetic waves interacting with targets. Therefore, channel models need to incorporate the electromagnetic scattering characteristics of different targets. Lastly, in modeling methods, it is necessary to consider the cascaded characteristics of channels with targets and the coupling with traditional communication channels[9, 10, 11, 12].
Currently, the standardization of ISAC channel models by 3GPP has reached an intermediate stage, defining ISAC scenarios, sensing targets, sensing modes, and channel model frameworks. The scenarios include traditional UMi, UMa, RMa, InH scenarios, as well as V2X, UAV, and industrial IoT scenarios. According to the scenarios, sensing targets are categorized into five types: humans, vehicles, UAVs, AGVs, and obstacles. Sensing modes are divided into Bi-Static and Mono-Static, and further subdivided into six sensing use cases based on whether the sensing device is a terminal or a base station: BS-UT-Bi, BS-BS-Bi, UT-UT-Bi, UT-BS-Mono, UT-UT-Mono, and BS-BS-Mono[13]. In the channel model framework, the ISAC channel is defined as the sensing channel plus the background channel. The sensing channel includes all parts of the channel interacting with targets, while the background channel includes all other parts of the channel that do not interact with targets. Based on the current progress of 3GPP standardization work and the GBSM simulation framework, this paper proposes an ISAC channel simulation framework aimed at future standardization and integrates it into the BUPTCMCC-6G-CMG+ channel simulation platform. The key characteristics of ISAC channels, such as cascaded characteristics, RCS polarization, scattering properties, and the coupling of target and background channels, are simulated and analyzed to provide a reference model system for ISAC channel simulation.
Figure 1: An ISAC channel paradigm
2 Simulation Architecture
This paper extends the 3GPP Geometry-Based Stochastic Model (GBSM) channel model by integrating new characteristics of sensing targets, thereby enhancing its scalability while maintaining the continuity of 5G channel models. This extension addresses the simulation requirements for evaluating the sensing performance of ISAC channels. We propose a novel ISAC channel model simulation architecture based on a cascaded channel approach. In this model, the ISAC channel is divided into two components: the target channel and the background channel. The geometric schematic of the model is shown in Figure 1.
The target channel is subdivided into Tx-Target and Target-Rx sub-channels.In the bi-static sensing mode, the target channel is formed by cascading the Tx-Target and Target-Rx sub-channels, with each sub-channel’s parameters generated using the 3GPP statistical model.In the mono-static sensing mode, the Tx-Target link is generated using the statistical model, while the Target-Rx part is derived utilizing channel reciprocity.
As for background channel in the bi-static sensing mode, the background channel is directly generated using the 3GPP statistical channel model.In the mono-static sensing mode, the background channel can either be evolved from the 3GPP method (where related work is limited) or extracted through additional measurements to obtain the channel model parameters.
It is crucial to consider the multipath power distribution characteristics between the target channel and the background channel. The power of both the target and background channels must be normalized to ensure that the overall channel conforms to the 3GPP loss parameters.
Table 1: 3GPP ISAC Standardization Focused Scenarios
            Sensing targets             Scenarios
            UAVs             RMa-AV UMa-AV UMi-AV
            Humans indoors             InF InH UMi(80%) UMa(80%) RMa(50%)
            Humans outdoors             UMi(20%) UMa(20%) RMa(50%)
            Automotive vehicles             Highway Urban grid
            Automated guided vehicles             InF
            Objects creating hazards on road/railways             Highway Urban grid HST
Figure 2: The steps for generating ISAC channel coefficients. (changes compared to the GBSM model are highlighted in gray.)
Our model framework is an extension of the 3GPP TR38.901 GBSM model[9]. Fig. 2 illustrates the overall simulation process. Users need to perform general configuration in step 1 (i.e., BS, UE, and target positions, antenna configuration, scene, target type). Steps 2-5 pertain to large-scale parameter generation, steps 6-12 focus on small-scale parameter generation, and steps 13-14 generate the channel coefficients.
In particular, steps 4, 9, and 12, which correspond to large-scale parameter cascading, RCS generation, and small-scale parameter cascading, are the extended steps to adapt to ISAC channel characteristics. It is important to note that the steps not highlighted in gray are consistent with the GBSM model generation process. However, each step needs to consider the channel parameter generation for the Tx-Target, Target-Rx, and background channels.
In step 4, the large-scale parameter cascading involves power normalization for the target and background channels. In step 12, the small-scale parameter cascading includes cluster reduction to decrease the simulation complexity.
2.1 Concatenate channel modeling
We propose a parameter cascading method for ISAC channel simulation based on the scenario layouts, path loss formulas, LOS probability, and large-scale parameters provided in the 3GPP channel model standards TR38.901, TR36.777, and TR37.885. Specifically, TR38.901 is utilized for human and AGV sensing target scenarios, TR36.777 for UAV target scenarios, and TR37.885 for V2X target scenarios. The proposed cascading simulation framework can be divided into large-scale parameter cascading and small-scale parameter cascading, which will be detailed in the following two subsections[14].
2.1.1 Large scale parameters concatenate
As illustrated in Fig. X, the simulation begins by setting up the simulation scenario, target type, target position, target movement speed and direction, number and location of base stations and terminals, antenna patterns and orientation angles of base stations and terminals, system center frequency, and terminal movement speed and direction. For the configured scenario, the LOS probability and path loss for the Tx-Target and Target-Rx links are calculated separately. Based on the target type, the corresponding RCS mean is generated, and combined with the path loss from both links, the cascaded path loss of the target channel can be written as
$PL_{target}(d_{1},d_{2},f_{0})=PL_{1}(d_{1},f_{0})+PL_{2}(d_{2},f_{0})+10\log{\frac{c^{2}}{4\pi f_{0}^{2}}}-10\log{\overline{\sigma}_{RCS}}$ (1)
Here, $d_{1}$ and $d_{2}$represents the distance from the transmitter to the target and the distance from the target to the receiver respectively; $c$ is the center frequency of the simulation,commonly taken as $3\times 10^{8}$m/s; $\overline{\sigma}_{RCS}$ denotes the mean radar cross-section area of the target.
2.1.2 Small scale parameters concatenate
The overall channel parameter generation expression used by the BUPTCMCC-6G-CMG+ channel simulator is as follows:
$H_{u,s,n}(t)=\sqrt{\frac{P_{n}}{M}}\sum_{m=1}^{M}\begin{bmatrix}F_{rx,u,\theta}(\theta_{n,m,ZOA},\phi_{n,m,AoA})\\ F_{rx,u,\phi}(\theta_{n,m,ZOA},\phi_{n,m,AoA})\end{bmatrix}^{T}\begin{bmatrix}exp(j\Phi_{n,m}^{\theta\theta})&\sqrt{\kappa_{n,m}^{-1}}exp(j\Phi_{n,m}^{\theta\phi})\\ \sqrt{\kappa_{n,m}^{-1}}exp(j\Phi_{n,m}^{\phi\theta})&exp(j\Phi_{n,m}^{\phi\phi})\end{bmatrix}$
$\begin{bmatrix}F_{tx,s,\theta}(\theta_{n,m,ZOD},\phi_{n,m,AoD})\\ F_{tx,s,\phi}(\theta_{n,m,ZOD},\phi_{n,m,AoD})\end{bmatrix}exp(j\frac{2\pi}{\lambda}\hat{r}_{rx,n,m}^{T}\cdot\overline{d}_{rx,u})exp(j\frac{2\pi}{\lambda}\hat{r}_{tx,n,m}^{T}\cdot\overline{d}_{tx,s})exp(j2\pi v_{n,m}t)$ (2)
Here, $P_{n}$ represents the power of the $n$-th cluster, generated in Step 7; $M$ denotes the number of sub-paths within each cluster; $F_{rx,u,\theta}$ and $F_{rx,u,\phi}$ are the field components of the receiving antenna in the $\theta_{n,m,ZOA}$ and $\phi_{n,m,AOA}$ directions, respectively, while $F_{tx,s,\theta}$ and $F_{tx,s,\phi}$ are the field components of the transmitting antenna in the $\theta_{n,m,ZOD}$ and $\phi_{n,m,AOD}$ directions, respectively. $\theta_{n,m}$ and $\phi_{n,m}$ are the vertical and horizontal angles of the $m$-th sub-path within the $n$-th cluster, generated in Step 8. $\hat{r}_{rx,n,m}$ and $\hat{r}_{tx,n,m}$ are the unit vectors in spherical coordinates, with the calculation expressions as follows:
$\hat{r}_{rx,n,m}=\begin{bmatrix}\sin{\theta_{n,m,ZOA}}\cos{\theta_{n,m,AOA}}\\ \sin{\theta_{n,m,ZOA}}\sin{\theta_{n,m,AOA}}\\ \cos{\theta_{n,m,ZOA}}\end{bmatrix},\hat{r}_{tx,n,m}=\begin{bmatrix}\sin{\theta_{n,m,ZOD}}\cos{\theta_{n,m,AOD}}\\ \sin{\theta_{n,m,ZOD}}\sin{\theta_{n,m,AOD}}\\ \cos{\theta_{n,m,ZOD}}\end{bmatrix}$ (3)
$\overline{d}_{rx,u}$ and $\overline{d}_{tx,s}$ represent the position vectors of the receiving and transmitting antenna elements, respectively; $\kappa_{n,m}$ is the linear value of the cross-polarization power ratio, generated in Step 10; $\lambda_{0}$ is the wavelength corresponding to the carrier frequency band. and the Doppler frequency factor $v_{n,m}$ is calculated as follows:
$v_{n,m}=v_{n,m}^{tx-tar}+v_{n,m}^{tar-rx}=\frac{\hat{r}_{tar,n,m}^{T}\cdot\overline{v}_{t}ar+\hat{r}_{rx,n,m}^{T}\cdot\overline{v}}{\lambda_{0}}$ (4)
where, $\overline{v}_{tar}$ and $\overline{v}$ represent the velocity vectors of the target and the receiver, respectively, and $\mathbf{r}$ is the unit vector in spherical coordinates at the target’s position can be expressed as:
$\hat{r}_{rx,n,m}=\begin{bmatrix}\sin{\theta_{n,m,ZOA}}\cos{\theta_{q,m,AOA}}\\ \sin{\theta_{q,m,ZOA}}\sin{\theta_{q,m,AOA}}\\ \cos{\theta_{q,m,ZOA}}\end{bmatrix}$ (5)
Considering that the cascaded link between the transmitter-target (tx-tar) and target-receiver (tar-rx) involves four possible scenarios as: $LOS+LOS$, $LOS+NLOS$, $NLOS+LOS$, and $NLOS+NLOS$. Equation 2.1.2 presents the general channel expression under the NLOS-NLOS condition. When LOS conditions are involved, the corresponding LOS components in Equation (1), such as power $P_{n}$, angles $\theta$ and $\phi$, and cross-polarization power ratio, should be selectively modified according to the expressions for the LOS case in the GBSM model. In step 4 of Fig. X, large-scale parameters are generated, producing delay, cluster power, angle, XPR, and other small-scale parameters for both the Tx-Target and Target-Rx links. The cascaded small-scale parameters are generated following the process outlined below. Clusters are pruned from the Tx-Target and Target-Rx links based on a predefined power threshold. Let $Q$ and $P$ be the number of clusters remaining in the Tx-Target and Target-Rx links, respectively, after pruning. The total number of clusters in the target channel is $N=QP$. For the n-th cluster of the target channel where $n=P$($q-1$)$+q$.
Cluster delay:
$\tau_{n}=\tau_{p}+\tau_{q}$ (6)
where, $\tau_{p}$ and $\tau_{q}$ represent the delay of p-th and q-th cluster from Tx-Target and Target-Rx links respectively.
Cluster power:
$P_{n}=P_{p}P_{q}$ (7)
Here, $P_{p}$ and $P_{q}$ represent the power of p-th and q-th cluster respectively.
Doppler frequency:
$exp(j2\pi\frac{v_{n}(t)}{\lambda_{0}})=exp(j2\pi\frac{v_{p}(t)}{\lambda_{0}})exp(j2\pi\frac{v_{q}(t)}{\lambda_{0}})$ (8)
where, $v_{p}(t)$ and $v_{q}(t)$ represent the doppler frequency of p-th and q-th links respectively; $\lambda_{0}$ donates the Wavelength corresponding to the center frequency.
Departure angle:
$\theta_{n,ZoD}=\theta_{q,ZoD},\theta_{n,AoD}=\theta_{q,AoD}$ (9)
Arrival angle:
$\theta_{n,ZoA}=\theta_{p,ZoA},\theta_{n,AoA}=\theta_{p,AoA}$ (10)
$\begin{bmatrix}exp(j\Phi_{n}^{\theta\theta})&\sqrt{\kappa_{n,m}^{-1}}exp(j\Phi_{n}^{\theta\phi})\\ \sqrt{\kappa_{n,m}^{-1}}exp(j\Phi_{n}^{\phi\theta})&exp(j\Phi_{n}^{\phi\phi})\end{bmatrix}=\begin{bmatrix}exp(j\Phi_{q}^{\theta\theta})&\sqrt{\kappa_{q}^{-1}}exp(j\Phi_{q}^{\theta\phi})\\ \sqrt{\kappa_{q}^{-1}}exp(j\Phi_{q}^{\phi\theta})&exp(j\Phi_{n}^{\phi\phi})\end{bmatrix}S\\ \begin{bmatrix}exp(j\Phi_{p}^{\theta\theta})&\sqrt{\kappa_{p}^{-1}}exp(j\Phi_{p}^{\theta\phi})\\ \sqrt{\kappa_{p}^{-1}}exp(j\Phi_{p}^{\phi\theta})&exp(j\Phi_{p}^{\phi\phi})\end{bmatrix}$ (11)
In the 3GPP RAN 1#116 meeting, RAN 1 defined the ISAC channel is composed of a component of target channel and a component of background channel, but how to generate the combined ISAC channel are considered for further study. The relationship between the component of background channel and the environment channel with no sensing target is still an open issue. we propose a new parameter named coupling factor to combine the path loss of target and background channels, and then the path loss of the ISAC channel can be expressed as:
$PL_{ISAC}=PL_{target}+O_{isac}\cdot PL_{backgraound}^{standardization}$ (12)
where the $PL_{ISAC}$ and $PL_{target}$ represent the linear path loss values of the ISAC channel and target channel, respectively. The $PL_{backgraound}^{standardization}$ represents the environment channel, i.e., the initial background channel with no sensing target, and it is important to note that $PL_{backgraound}^{standardization}$ is not the same as the defined $PL_{backgraound}$he latter represents the path loss of the channel components after removing the target channel components. Considering the similarity and compatibility with the communication channel, it is suggested to reuse the 3GPP TR 38.901 path loss. The values of $O_{isac}$ needs for further study with more measurement results.
2.2 RCS modeling
Radar utilizes the scattered power of targets for sensing, and the modeling of the Radar Cross Section (RCS) of targets directly impacts the performance evaluation of sensing. First, the concept of the ”three-dimensional resolution cell” needs to be introduced. A resolution cell is defined as the volume instantaneously illuminated and scattered, comprising two dimensions: angle and distance, and is defined as follows:
$V=\frac{\Omega R^{2}c\tau}{2}$ (13)
where $\frac{c\tau}{2}$ represents the distance resolution cell, $c$ is the speed of light, and $\tau$ is the pulse width; $\Omega$ donates be the antenna beam solid angle, and $R$ is the distance from the radar receiver to the target. If the target’s size is contained within the volume $V$, the target is considered a point target. For the five types of sensing targets defined by 3GPP UAV, Human, Vehicle at a long distance, AGV, and small road obstacles—modeling as point targets is appropriate. For targets larger than the resolution cell and with irregular shapes, such as vehicles and AGVs at close range, large obstacles, and humans in gesture recognition scenarios, multi-point RCS modeling may be necessary.
Table 1 provides sample RCS values (average values) for common targets in the microwave frequency band.
2.2.1 RCS polarization
The scattering characteristics of the target are related to the polarization of the incident field. When the antenna is linearly polarized, the linearly polarized wave incident on the target surface can be decomposed into vertical and horizontal polarization components. The polarization components of the target’s scattered field received by the receiving antenna can be expressed as:
$\begin{bmatrix}E_{H}^{R}\\ E_{V}^{R}\end{bmatrix}=\begin{bmatrix}\alpha_{HH}&\alpha_{VH}\\ \alpha_{HV}&\alpha_{VV}\end{bmatrix}\begin{bmatrix}E_{H}^{T}\\ E_{V}^{T}\end{bmatrix}$ (14)
where $E_{H}^{R}$ and $E_{V}^{R}$denote the horizontally and vertically polarized components of the received electric field, respectively, $E_{H}^{T}$ and $E_{V}^{T}$ denote the horizontally and vertically polarized components of the transmitted electric field, respectively. The scattering matrix $S$ represents the coefficients related to the target’s scattering characteristics and polarization.
2.2.2 Swerling modeling
For complex targets, when the spacing between scattering elements is comparable to the wavelength, changes in the observation direction or operating wavelength alter the phase of each scattering channel received at the input of the receiver, thereby modifying the resultant vector and producing fluctuating target signals. The size of the Radar Cross Section (RCS) directly affects radar detection performance, and in engineering calculations, RCS is often treated as a constant, as shown by the average values in Table 1. For targets in motion, the cross-sectional area fluctuates with the viewing angle. To accurately characterize the fluctuations of the radar cross section, its probability density function $p(\sigma)$ must be provided. The Swerling fluctuation model is a classic and widely used model in radar applications. It categorizes fluctuations into four types based on the correlation time and fluctuation rate:
Swerling 1 and 2:
$p(\sigma)=\frac{1}{\overline{\sigma}}exp(-\frac{\sigma}{\overline{\sigma}})$ (15)
where $\overline{\sigma}$ is the average value of RCS, $\sigma$ follows an exponential distribution and $\sigma>0$.The probability distribution functions of Swerling 1 and Swerling 2 are Identical, but Swerling 1 Exhibits Slow Fluctuations While Swerling 2 Exhibits Fast Fluctuations.
Swerling 3 and 4:
$p(\sigma)=\frac{4\sigma}{\overline{\sigma}^{2}}exp(-\frac{2\sigma}{\overline{\sigma}})$ (16)
The probability distribution functions of Swerling 3 and Swerling 4 are Identical, in which Swerling 3 Exhibits Slow Fluctuations and Swerling 4 Exhibits Fast Fluctuations.The first and second types of fluctuation models are applicable to complex targets composed of a large number of approximately equal scatterers, such as an aircraft.The third and fourth types of fluctuation models are suitable for targets that consist of one large reflecting surface combined with many small scattering surfaces or for a single large scattering surface with small changes in aspect.
2.2.3 Measurement based modeling
Based on the extensive frequency, distance, and target type-dependent measurement analysis conducted by the team in [15], the following conclusions can be drawn regarding RCS modeling:
  • The RCS of the UAV is angle-independent and exhibits similar properties to omnidirectional scattering. In addition, the measurement results from mono-static mode and bi-static mode are found to be closely approximated.
  • The RCS model of the UAV is modeled as a normal distribution in both mono-static and bi-static sensing modes. The mean is used to determine the deterministic component and the variance is used to determine the randomly generated component.
  • The RCS of the human is angle-independent and exhibits similar properties to omnidirectional scattering. In addition, the measurement results from mono-static mode and bi-static mode are found to be closely approximated.
  • The RCS model of the human is modeled as a normal distribution in both mono-static and bi-static sensing modes. The mean is used to determine the deterministic component and the variance is used to determine the randomly generated component.
Different from human body and UAV, vehicle RCS have spiking characteristics and are not suitable for fitting with omnidirectional patterns. The RCS of humans and UAVs can be fitted using a log-normal distribution, whereby the mean $\mu$ and variance $\sigma$ are derived from empirical measurements for random generation. The RCS of vehicles exhibits four peaks in its angular distribution, necessitating further discussion on its modeling and simulation methods.
$RCS_{Bi}=RCS_{Mono}\cos{\frac{\beta}{2}}$ (17)
This equation holds under the condition that the target is sufficiently smooth, without the presence of shadowing and backscatter. The bistatic RCS for a specific target requires support from measurement results and can be modeled similarly to an antenna radiation pattern.
3 Evaluation Metrics for ISAC Testing
3.1 Accuracy metric
Table 2: 3GPP ISAC Standardization Focused Scenarios
         Sensing usage          Principle          Accuracy          Resolution
         Ranging          $R=\frac{1}{2}ct_{R}$          $dR=\frac{1}{2}cdt_{R}+\frac{R}{C}dc$          $\Delta R=\frac{1}{2}ct_{\tau}$
         Speed detection          $f_{d}=\frac{2v}{\lambda}$          $dv=\frac{\lambda}{2}df_{d}$          $\Delta v=\frac{\lambda}{2}\Delta f_{d}$
         Angulation          $d\theta=\arcsin(\frac{\phi\lambda}{2\pi d})$          $d\theta=\frac{\lambda d\phi}{2\pi d\cos{\theta}}$          $\Delta\theta>\frac{\lambda}{Nd\cos{\theta}}$
The basic detection equation in the field of radar is as follows in 2. Here,$c$ is the speed of light with a value of $3\times 10^{8}m/s$ and $t_{R}$ represents the time difference between the transmitted pulse and the received echo pulse from the sensed target. The ranging accuracy can be qualitatively analyzed using the total differential expression of equation LABEL:eq_RDtheory. The range resolution $\Delta R$ is defined as the minimum distinguishable distance, which is related to the pulse width $t_{\tau}$ of the sensing signal. The maximum unambiguous range is determined by the pulse repetition interval $T_{R}$ and is expressed as follows:
$R_{Max}=\frac{1}{2}cT_{R}$ (18)
$R=\frac{1}{2}c(mT_{R}+t_{r})$ (19)
The basic principle of angle measurement in the radar field is as follows.
$\phi=\frac{2\pi}{\lambda}d\sin{\theta}$ (20)
where $\phi$ represents the phase difference between different antenna elements that are separated by a distance $d$ at the receiver and $\theta$ is the angle of sensing target.The angle measurement accuracy can be expressed as
$d\phi=\frac{2\pi}{\lambda}d\cos{\theta}d\theta$ (21)
$\phi=\frac{2\pi}{\lambda}d\sin{\theta}<\pi$ (22)
from equation 22,$\theta_{max}=\arcsin{(\frac{\lambda}{2\pi})}$. The angle resolution $\Delta\theta$ is defined as the minimum distinguishable angle, which is related to the main lobe width of the antenna array $\frac{2\pi}{N}$. according to LABEL:eq_ADresolution,$\Delta\theta>\frac{\lambda}{Nd\cos{\theta}}$ where $N$ is the number of antenna array. the basic principle of velocity measurement is based on the Doppler frequency. from the above equation,we obtain $v=\frac{\lambda f_{d}}{2}$.analogous to range resolution and angular resolution. Here, the Doppler resolution $\Delta f_{d}=\frac{1}{MT_{R}}$, in which $M$ is the number of transmitted pulses, and $T_{R}$ is the pulse repetition interval.
3.2 Detection metric
False alarm probability occurs when the noise at the receiver exceeds the decision threshold, leading to a false detection. This probability is related to the root mean square (RMS) value of the noise and the decision threshold, and is defined as follows.
[…]
where $V_{T}$ is the decision threshold, $\omega(x)$ denotes the probability density function (PDF) of detection.When the noise is complex and both its real and imaginary components follow a Gaussian distribution, the magnitude of the noise follows a Rayleigh distribution,$\omega(x)$ can be expressed as
$\omega(x|\text{no target})=\frac{x}{\sigma^{2}}exp(-\frac{x^{2}}{2\sigma^{2}}),x\geq 0$ (24)
substituting Equation 24 into Equation 23 results in
$P_{fa}=exp(-\frac{V_{T}^{2}}{2\sigma^{2}})$ (25)
further, the relationship between $P_{fa}$ and $V_{T}$ is obtained as $V_{T}=\sqrt{2\sigma\ln{P_{fa}}}$ The probability of detection, $P_{D}$, represents the likelihood of correctly detecting a target when one is present and is defined as follows.
[…]
4 Simulation Results
In the simulation experiments, the focus is primarily on the reliability of the ISAC channel cascading method and the RCS simulation results. To validate the channel cascading method, specific scenario parameters used in the dual-station sensing mode of a UAV scenario are presented in Table X. The statistical characteristics of the generated channels, including delay spread, horizontal angle of arrival (AoA) spread, horizontal angle of departure (AoD) spread, vertical AoA spread, and vertical AoD spread, are validated for the 5G standard channel model, full convolution method, and the proposed parameter convolution method. The number of simulation runs is set to 1000 to collect statistics on the generated channel’s spread values. For the validation of the RCS simulation results, the mean and variance of the RCS concerning the log-normal distribution are obtained by fitting the measured RCS results reported by the team in document [15]. The RCS is simulated based on the fitted distribution, and the simulated RCS results are compared against the measured data to evaluate the error.
Figure 3: Delay sperad of ISAC concatenated Channel
As shown in Fig. 3, the delay spread of the cascaded channels is always greater than that of the standard channel, which is attributed to the addition of the two delay components during the cascading process, resulting in a broader overall delay distribution function and an increase in the spread value statistics. The spread value under the parameter convolution method is smaller than that of the full convolution method, as the parameter convolution approach reduces complexity through cluster pruning, thereby eliminating low-power, high-delay clusters, which in turn reduces the spread value.
ASA of ISAC Channel
ASD of ISAC Channel
ESA of ISAC Channel
ESD of ISAC Channel
Figure 4: Angle spread of ISAC Channel
RCS Human 5GHz
RCS Human 14GHz
RCS Human 28GHz
RCS UAV 5GHz
RCS UAV 14GHz
RCS UAV 28GHz
Figure 5: RCS of UAV and Human in 5GHz 14GHz 28GHz respectively
Fig. 4 illustrates that the statistical values of the angle spread for the three channels closely align, especially in the horizontal angle spread dimension. This is due to the cascading scheme generating statistical parameters for each segment, while the mean and variance distributions remain consistent across simulations of the same scenario, leading to a convergence of spread values through multiple statistical averages. A deviation is observed in the ESD, where the spread value of the cascaded scheme exceeds that of the standard channel. This is because the cluster pruning retains high-power clusters, which are closely aligned with the angle of arrival. Consequently, after pruning and cascading, high-power clusters are duplicated, leading to an increase in the spread value. The spread value of the parameter convolution method is lower than that of the full convolution due to the reduction of clusters with larger angles after pruning. Fig. 5 presents the RCS simulation results for two target types, Human and UAV, at frequencies of 5 GHz, 14 GHz, and 28 GHz. The simulated results align well with the measured results reported in proposal x, validating the feasibility of the log-normal fitting scheme. Additionally, it confirms the differences in RCS distribution between integrated sensing and communication channels and traditional radar systems. This discrepancy arises because traditional radar RCS perception occurs at longer distances, whereas integrated sensing and communication services focus on closer ranges, leading to variations between the RCS distributions in radar applications and the measured RCS results in ISAC.
5 Conclusion
In this paper, we extended the 3D-GBSM 5G standard channel model, integrating the latest advancements in 3GPP ISAC standardization to enhance the sensing and communication integrated channel simulation capabilities for 6G. The new model supports both Bi-Static and Mono-Static sensing modes, accommodating four types of sensing targets: UAVs, vehicles, humans, and AGVs. It aligns with various options outlined in the 3GPP standardization progress, including modeling methods for RCS size and scale, polarization techniques, and the power normalization methods for target and background channels.
By employing a parameter cascading method, we preserved the statistical characteristics of the ISAC channel distribution while achieving a 93.75% reduction in average convolution complexity. The simulation database for target RCS was developed through the fitting of actual measurements related to corresponding target RCS. A comparative analysis demonstrated good consistency between the simulated RCS and the measured RCS.
This advancement accelerates the standardization of the 6G ISAC channel model and the evaluation of new algorithms, as we can now utilize existing 5G standard parameter tables along with the proposed cascading model and measured RCS data to simulate ISAC channels. The modified simulation framework is applicable to all future developments in ISAC channel model standardization, providing insights for the standardization simulation process and expediting simulation evaluation progress.
[…]
Appendix A.