### v2

# ISAC channel simulator GBSM

Abstract
Integrated Sensing and Communication (ISAC) represents a pivotal technology in 6G, with ongoing research and standardization efforts advancing robustly. This paper presents a design and implementation methodology for ISAC channel simulation, utilizing a Geometry-Based Stochastic Model (GBSM) framework. First, we introduce the progress of 3GPP ISAC channel standardization and the key topics of discussion. Second, in response to the existing gap in standardized ISAC channel simulation frameworks, we propose a cascaded ISAC channel simulation framework founded on the Geometry-Based Stochastic Model (GBSM), utilizing our team’s relevant measurements, analyses, and proposed results. Building upon this framework, we have developed and designed the ISAC channel simulator, BUPTCMCC-6G-CMG+. Finally, we analyze and validate the results of the simulation platform, and present prospects for future ISAC testing research in conjunction with channel simulators.
keywords:
Channel simulator, GBSM, concatenate, ISAC testing, ISAC standardization.
1 Introduction
Next-generation mobile communication technology (6G) is anticipated to enhance traditional communication capabilities, including high capacity, low latency, and high reliability, while also achieving native intelligence, ISAC, and ubiquitous connectivity. The boundary between the physical and digital realms is expected to become increasingly indistinct, thereby offering users a ubiquitous and immersive experience [1, 2]. ISAC technology is expected to facilitate the implementation of diverse use cases, particularly in target detection, localization, tracking, imaging, and the estimation of target type, speed, and angle. These capabilities will enable applications in autonomous driving, industrial automation, digital twin reconstruction of physical environments, gesture detection, and activity recognition. The sensing information derived from ISAC can facilitate the development of smart homes, factories, smart cities, and intelligent transportation systems. Additionally, ISAC technology introduces new performance dimensions relative to traditional communication technologies, including detection probability, speed, angle, delay, sensing accuracy, resolution, and sensing capacity density[6].
The transmission medium at both the transmitter and receiver establishes the performance limits of wireless communication systems. Furthermore, the wireless channel serves as a critical component in the research and evaluation of wireless communication technologies[7]. The evaluation scenarios for ISAC introduce new key requirements for channel models, particularly regarding sensing modes, target modeling, and modeling methodologies. Firstly, regarding sensing modes, traditional channel models typically represent bi-static channels used for communication purposes. Sensing scenarios necessitate the use of mono-static channel models for self-sensing; therefore, ISAC channel models must incorporate both mono-static and bi-static sensing configurations. Secondly, in the context of target modeling for sensing, ISAC applications expand the evaluation criteria to include target detection, localization, identification, and tracking. Examples include intrusion detection, tracking of UAVs and V2X communications, as well as recognition of various types and sizes of targets. This necessitates the utilization of scattering properties, polarization characteristics, and Doppler information of electromagnetic waves as they interact with targets. Consequently, channel models must integrate the electromagnetic scattering characteristics of various targets. Finally, in modeling methods, it is imperative to account for the cascaded characteristics of channels interacting with targets, as well as their coupling with traditional communication channels.[9, 10, 11, 12].
At present, the standardization of ISAC channel models by 3GPP has reached an intermediate stage, wherein ISAC scenarios, sensing targets, sensing modes, and channel model frameworks have been defined. The defined scenarios encompass traditional Urban Micro (UMi), Urban Macro (UMa), Rural Macro (RMa), and Indoor Hotspot (InH) scenarios, in addition to Vehicle-to-Everything (V2X), Unmanned Aerial Vehicle (UAV), and Industrial Internet of Things (IoT) scenarios. Based on the defined scenarios, sensing targets are classified into five categories: humans, vehicles, unmanned aerial vehicles (UAVs), automated guided vehicles (AGVs), and obstacles. Sensing modes are categorized into bi-static and mono-static configurations, which are further subdivided into six specific use cases based on whether the sensing device is a user terminal (UT) or a base station (BS): BS-UT-Bi, BS-BS-Bi, UT-UT-Bi, UT-BS-Mono, UT-UT-Mono, and BS-BS-Mono.[13]. In the framework of channel modeling, the ISAC channel is defined as the combination of the sensing channel and the background channel. The sensing channel encompasses all components of the channel that interact with targets, whereas the background channel comprises all other components that do not engage with targets.Building upon the current advancements in 3GPP standardization efforts and the GBSM simulation framework, this paper proposes an ISAC channel simulation framework designed for future standardization, which is integrated into the BUPTCMCC-6G-CMG+ channel simulation platform. The key characteristics of ISAC channels—including cascaded features, RCS polarization, scattering properties, and the coupling between target and background channels—are simulated and analyzed to establish a reference model system for ISAC channel simulation.
Figure 1: An ISAC channel paradigm
2 Simulation Architecture
This paper expands the 3GPP Geometry-Based Stochastic Model (GBSM) channel model by incorporating new characteristics of sensing targets, thereby improving its scalability while preserving the continuity of 5G channel models. This expansion fulfills the simulation requirements for assessing the sensing performance of ISAC channels. We propose an innovative ISAC channel model simulation architecture grounded in a cascaded channel framework. In this model, the ISAC channel is classified into two components: the target channel and the background channel. The geometric schematic of the model is shown in 1.
The target channel is further classified into Tx-Target and Target-Rx sub-channels. In the bi-static sensing mode, the target channel is constructed by cascading the Tx-Target and Target-Rx sub-channels, with parameters for each sub-channel generated according to the 3GPP statistical model. In the mono-static sensing mode, the Tx-Target link is generated using the statistical model, whereas the Target-Rx component is derived by utilizing channel reciprocity.
In the bi-static sensing mode, the background channel is directly generated using the 3GPP statistical channel model. In the mono-static sensing mode, the background channel can be derived from the 3GPP method, which has limited related work, or extracted through additional measurements to obtain the necessary channel model parameters.
Considering the multipath power distribution characteristics between the target channel and the background channel is essential. Both the target and background channel powers must be normalized to ensure that the overall channel adheres to the 3GPP loss parameters.
Table 1: 3GPP ISAC Standardization Focused Scenarios
            Sensing targets             Scenarios
            UAVs             RMa-AV UMa-AV UMi-AV
            Humans indoors             InF InH UMi(80%) UMa(80%) RMa(50%)
            Humans outdoors             UMi(20%) UMa(20%) RMa(50%)
            Automotive vehicles             Highway Urban grid
            Automated guided vehicles             InF
            Objects creating hazards on road/railways             Highway Urban grid HST
Figure 2: The steps for generating ISAC channel coefficients. (changes compared to the GBSM model are highlighted in gray.)
Our model framework extends the 3GPP TR38.901 GBSM model[9]. Fig. 2 depicts the overall simulation process. In Step 1, users are required to perform general configurations, including the positions of the base station (BS), user equipment (UE), and targets, as well as antenna configuration, scene settings, and target types. Steps 2 to 5 are dedicated to large-scale parameter generation, while Steps 6 to 12 concentrate on small-scale parameter generation. Finally, Steps 13 and 14 are responsible for generating the channel coefficients.
Specifically, Steps 4, 9, and 12, which correspond to large-scale parameter cascading, radar cross-section (RCS) generation, and small-scale parameter cascading, are extended to accommodate the characteristics of ISAC channels. It is essential to emphasize that the steps not highlighted in gray align with the GBSM model generation process. However, each step must take into account the generation of channel parameters for the Tx-Target, Target-Rx, and background channels.
In step 4, the cascading of large-scale parameters entails the normalization of power for both the target and background channels. In step 12, the cascading of small-scale parameters incorporates cluster reduction to mitigate simulation complexity.
2.1 Concatenate channel modeling
We propose a parameter cascading method for ISAC channel simulation, grounded in the scenario layouts, path loss formulas, line-of-sight (LOS) probabilities, and large-scale parameters delineated in the 3GPP channel model standards TR38.901, TR36.777, and TR37.885. Specifically, TR38.901 is employed for scenarios involving human and automated guided vehicle (AGV) sensing targets, TR36.777 is designated for unmanned aerial vehicle (UAV) target scenarios, and TR37.885 is applied to vehicle-to-everything (V2X) target scenarios. The proposed cascading simulation framework comprises two main components: large-scale parameter cascading and small-scale parameter cascading, which will be elaborated upon in the subsequent two subsections.[14].
2.1.1 Large scale parameters concatenate
As illustrated in Fig.2, the simulation commences with the establishment of various parameters, including the simulation scenario, target type, target position, movement speed and direction of the target, the number and location of base stations and terminals, antenna patterns, orientation angles of both base stations and terminals, system center frequency, and movement speed and direction of the terminals. For the configured scenario, the Line-of-Sight (LOS) probability and path loss for the Tx-Target and Target-Rx links are computed individually. Based on the target type, the corresponding Radar Cross Section (RCS) mean is generated. This value, when combined with the path loss from both links, allows the cascaded path loss of the target channel to be expressed as
$PL_{target}(d_{1},d_{2},f_{0})=PL_{1}(d_{1},f_{0})+PL_{2}(d_{2},f_{0})+10\log{\frac{c^{2}}{4\pi f_{0}^{2}}}-10\log{\overline{\sigma}_{RCS}}$ (1)
In this context, $d_{1}$ and $d_{2}$ represent the distances from the transmitter to the target and from the target to the receiver, respectively. Here, $c$ denotes the center frequency of the simulation, typically considered as $3\times 10^{8}$ m/s, while $\overline{\sigma}_{RCS}$ indicates the mean radar cross-section area of the target.
2.1.2 Small scale parameters concatenate
The overall expression for channel parameter generation utilized by the BUPTCMCC-6G-CMG+ channel simulator is as follows:
$H_{u,s,n}(t)=\sqrt{\frac{P_{n}}{M}}\sum_{m=1}^{M}\begin{bmatrix}F_{rx,u,\theta}(\theta_{n,m,ZOA},\phi_{n,m,AoA})\\ F_{rx,u,\phi}(\theta_{n,m,ZOA},\phi_{n,m,AoA})\end{bmatrix}^{T}\begin{bmatrix}exp(j\Phi_{n,m}^{\theta\theta})&\sqrt{\kappa_{n,m}^{-1}}exp(j\Phi_{n,m}^{\theta\phi})\\ \sqrt{\kappa_{n,m}^{-1}}exp(j\Phi_{n,m}^{\phi\theta})&exp(j\Phi_{n,m}^{\phi\phi})\end{bmatrix}$
$\begin{bmatrix}F_{tx,s,\theta}(\theta_{n,m,ZOD},\phi_{n,m,AoD})\\ F_{tx,s,\phi}(\theta_{n,m,ZOD},\phi_{n,m,AoD})\end{bmatrix}exp(j\frac{2\pi}{\lambda}\hat{r}_{rx,n,m}^{T}\cdot\overline{d}_{rx,u})exp(j\frac{2\pi}{\lambda}\hat{r}_{tx,n,m}^{T}\cdot\overline{d}_{tx,s})exp(j2\pi v_{n,m}t)$ (2)
Here, $P_{n}$ represents the power of the $n$-th cluster, generated in Step 7; $M$ denotes the number of sub-paths within each cluster; $F_{rx,u,\theta}$ and $F_{rx,u,\phi}$ are the field components of the receiving antenna in the $\theta_{n,m,ZOA}$ and $\phi_{n,m,AOA}$ directions, respectively, while $F_{tx,s,\theta}$ and $F_{tx,s,\phi}$ are the field components of the transmitting antenna in the $\theta_{n,m,ZOD}$ and $\phi_{n,m,AOD}$ directions, respectively. $\theta_{n,m}$ and $\phi_{n,m}$ are the vertical and horizontal angles of the $m$-th sub-path within the $n$-th cluster, generated in Step 8. $\hat{r}_{rx,n,m}$ and $\hat{r}_{tx,n,m}$ are the unit vectors in spherical coordinates, with the calculation expressions as follows:
$\hat{r}_{rx,n,m}=\begin{bmatrix}\sin{\theta_{n,m,ZOA}}\cos{\theta_{n,m,AOA}}\\ \sin{\theta_{n,m,ZOA}}\sin{\theta_{n,m,AOA}}\\ \cos{\theta_{n,m,ZOA}}\end{bmatrix},\hat{r}_{tx,n,m}=\begin{bmatrix}\sin{\theta_{n,m,ZOD}}\cos{\theta_{n,m,AOD}}\\ \sin{\theta_{n,m,ZOD}}\sin{\theta_{n,m,AOD}}\\ \cos{\theta_{n,m,ZOD}}\end{bmatrix}$ (3)
$\overline{d}_{rx,u}$ and $\overline{d}_{tx,s}$ represent the position vectors of the receiving and transmitting antenna elements, respectively; $\kappa_{n,m}$ is the linear value of the cross-polarization power ratio, generated in Step 10; $\lambda_{0}$ is the wavelength corresponding to the carrier frequency band. and the Doppler frequency factor $v_{n,m}$ is calculated as follows:
$v_{n,m}=v_{n,m}^{tx-tar}+v_{n,m}^{tar-rx}=\frac{\hat{r}_{tar,n,m}^{T}\cdot\overline{v}_{t}ar+\hat{r}_{rx,n,m}^{T}\cdot\overline{v}}{\lambda_{0}}$ (4)
where, $\overline{v}_{tar}$ and $\overline{v}$ represent the velocity vectors of the target and the receiver, respectively, and $\mathbf{r}$ is the unit vector in spherical coordinates at the target’s position can be expressed as:
$\hat{r}_{rx,n,m}=\begin{bmatrix}\sin{\theta_{n,m,ZOA}}\cos{\theta_{q,m,AOA}}\\ \sin{\theta_{q,m,ZOA}}\sin{\theta_{q,m,AOA}}\\ \cos{\theta_{q,m,ZOA}}\end{bmatrix}$ (5)
Considering that the cascaded link between the transmitter-target (tx-tar) and target-receiver (tar-rx) involves four possible scenarios: $LOS+LOS$, $LOS+NLOS$, $NLOS+LOS$, and $NLOS+NLOS$. Equation 2.1.2 presents the general channel expression under the NLOS-NLOS condition. When LOS conditions are involved, the corresponding LOS components in Equation (1), such as power $P_{n}$, angles $\theta$ and $\phi$, and cross-polarization power ratio, should be selectively modified according to the expressions for the LOS case in the GBSM model. In Step 4 of Fig. 2, Large-scale parameters are generated, resulting in the production of delay, cluster power, angles, XPR, and other small-scale parameters for both the Tx-Target and Target-Rx links. Subsequently, the cascaded small-scale parameters are generated in accordance with the outlined process. Clusters are pruned from the Tx-Target and Target-Rx links based on a predefined power threshold. Let $Q$ and $P$ be the number of clusters remaining in the Tx-Target and Target-Rx links, respectively, after pruning. The total number of clusters in the target channel is $N=QP$. For the n-th cluster of the target channel where $n=P$($q-1$)$+q$.
Cluster delay:
$\tau_{n}=\tau_{p}+\tau_{q}$ (6)
where $\tau_{p}$ and $\tau_{q}$ represent the delay of the p-th and q-th cluster from Tx-Target and Target-Rx links respectively.
Cluster power:
$P_{n}=P_{p}P_{q}$ (7)
Here, $P_{p}$ and $P_{q}$ represent the power of the p-th and q-th cluster respectively.
Doppler frequency:
$exp(j2\pi\frac{v_{n}(t)}{\lambda_{0}})=exp(j2\pi\frac{v_{p}(t)}{\lambda_{0}})exp(j2\pi\frac{v_{q}(t)}{\lambda_{0}})$ (8)
where, $v_{p}(t)$ and $v_{q}(t)$ represent the doppler frequency of p-th and q-th links respectively; $\lambda_{0}$ donates the Wavelength corresponding to the center frequency.
Departure angle:
$\theta_{n,ZoD}=\theta_{q,ZoD},\theta_{n,AoD}=\theta_{q,AoD}$ (9)
Arrival angle:
$\theta_{n,ZoA}=\theta_{p,ZoA},\theta_{n,AoA}=\theta_{p,AoA}$ (10)
Cross polarization power ratio (XPR):
$XPR=\begin{bmatrix}exp(j\Phi_{q}^{\theta\theta})&\sqrt{\kappa_{q}^{-1}}exp(j\Phi_{q}^{\theta\phi})\\ \sqrt{\kappa_{q}^{-1}}exp(j\Phi_{q}^{\phi\theta})&exp(j\Phi_{n}^{\phi\phi})\end{bmatrix}S\\ \begin{bmatrix}exp(j\Phi_{p}^{\theta\theta})&\sqrt{\kappa_{p}^{-1}}exp(j\Phi_{p}^{\theta\phi})\\ \sqrt{\kappa_{p}^{-1}}exp(j\Phi_{p}^{\phi\theta})&exp(j\Phi_{p}^{\phi\phi})\end{bmatrix}$ (11)
During the 3GPP RAN 1 #116 meeting, RAN 1 defined the ISAC channel as comprising a component from the target channel and a component from the background channel. However, the methodology for generating the combined ISAC channel is subject to further study. The relationship between the background channel component and the environment channel in the absence of sensing targets remains an unresolved issue. To address this, we introduce a new parameter, termed the coupling factor, to integrate the path loss of both the target and background channels. Consequently, the path loss of the ISAC channel can be expressed as follows:
$PL_{ISAC}=PL_{target}+O_{isac}\cdot PL_{backgraound}^{standardization}$ (12)
where the $PL_{ISAC}$ and $PL_{target}$ represent the linear path loss values of the ISAC channel and target channel, respectively. The $PL_{background}^{standardization}$ represents the environment channel, i.e., the initial background channel with no sensing target, and it is important to note that $PL_{background}^{standardization}$ is not the same as the defined $PL_{backgraound}$ The latter refers to the path loss of the channel components after the removal of the target channel components. Given the similarities and compatibility with the communication channel, it is advisable to reuse the path loss model from 3GPP TR 38.901. The values of $O_{isac}$ need for further study with more measurement results.
2.2 RCS modeling
Radar systems leverage the scattered power of targets for sensing, and the accurate modeling of the Radar Cross Section (RCS) of these targets directly influences the performance evaluation of sensing capabilities. First, it is essential to introduce the concept of the ”three-dimensional resolution cell.” A resolution cell is defined as the volume that is instantaneously illuminated and scattered, characterized by two dimensions: angle and distance. It can be formally expressed as follows:
$V=\frac{\Omega R^{2}c\tau}{2}$ (13)
where $\frac{c\tau}{2}$ represents the distance resolution cell, $c$ is the speed of light, and $\tau$ is the pulse width; $\Omega$ donates be the antenna beam solid angle, and $R$ is the distance from the radar receiver to the target. If the target’s size is contained within the volume $V$, the target is considered a point target. For the five types of sensing targets defined by 3GPP—UAVs, humans, long-distance vehicles, AGVs, and small road obstacles—modeling as point targets is appropriate. However, for targets that exceed the resolution cell and possess irregular shapes, such as close-range vehicles and AGVs, large obstacles, and humans in gesture recognition scenarios, multi-point RCS modeling may be required.
2.2.1 RCS polarization
The scattering characteristics of a target are influenced by the polarization of the incident field. When the antenna is linearly polarized, the incident linearly polarized wave on the target surface can be decomposed into vertical and horizontal polarization components. The polarization components of the scattered field from the target, as received by the receiving antenna, can be expressed as:
$\begin{bmatrix}E_{H}^{R}\\ E_{V}^{R}\end{bmatrix}=\begin{bmatrix}\alpha_{HH}&\alpha_{VH}\\ \alpha_{HV}&\alpha_{VV}\end{bmatrix}\begin{bmatrix}E_{H}^{T}\\ E_{V}^{T}\end{bmatrix}$ (14)
where $E_{H}^{R}$ and $E_{V}^{R}$denote the horizontally and vertically polarized components of the received electric field, respectively, $E_{H}^{T}$ and $E_{V}^{T}$ denote the horizontally and vertically polarized components of the transmitted electric field, respectively. The scattering matrix $S$ represents the coefficients related to the target’s scattering characteristics and polarization.
2.2.2 Swerling modeling
For complex targets, when the spacing between scattering elements is comparable to the wavelength, changes in the observation direction or operating wavelength alter the phase of each scattering channel received at the input of the receiver, thereby modifying the resultant vector and producing fluctuating target signals. The size of the Radar Cross Section (RCS) directly influences radar detection performance. In engineering calculations, RCS is frequently treated as a constant, typically represented by its average values. For moving targets, the RCS varies with the viewing angle. To accurately characterize these fluctuations, it is essential to provide its probability density function $p(\sigma)$. The Swerling fluctuation model is a classic and widely utilized framework in radar applications. It categorizes fluctuations into four distinct types based on correlation time and fluctuation rate:
Swerling 1 and 2:
$p(\sigma)=\frac{1}{\overline{\sigma}}exp(-\frac{\sigma}{\overline{\sigma}})$ (15)
where $\overline{\sigma}$ is the average value of RCS, $\sigma$ follows an exponential distribution and $\sigma>0$.The probability distribution functions of Swerling 1 and Swerling 2 are identical; however, Swerling 1 exhibits slow fluctuations, while Swerling 2 demonstrates fast fluctuations.
Swerling 3 and 4:
$p(\sigma)=\frac{4\sigma}{\overline{\sigma}^{2}}exp(-\frac{2\sigma}{\overline{\sigma}})$ (16)
The probability distribution functions of Swerling 3 and Swerling 4 are identical; however, Swerling 3 exhibits slow fluctuations, while Swerling 4 demonstrates fast fluctuations. The first and second types of fluctuation models apply to complex targets composed of a large number of approximately equal scatterers, such as aircraft. The third and fourth types of fluctuation models are appropriate for targets characterized by a large reflecting surface in conjunction with numerous small scattering surfaces, or for a singular large scattering surface exhibiting minor variations in aspect.
2.2.3 Measurement based modeling
Drawing upon the comprehensive analysis of measurements that depend on frequency, distance, and target type conducted by the team in [15], the following conclusions regarding RCS modeling can be established:
  • The RCS of the UAV is independent of angle and demonstrates characteristics akin to omnidirectional scattering. Furthermore, the measurement results obtained from both mono-static and bi-static modes closely approximate each other.
  • The RCS model for the UAV is represented as a normal distribution in both mono-static and bi-static sensing modes. The mean defines the deterministic component, while the variance characterizes the stochastic component.
  • The RCS of a human is angle-independent and displays characteristics akin to omnidirectional scattering. Furthermore, the measurement results obtained from mono-static and bi-static modes closely align with each other.
  • The RCS model for humans is represented as a normal distribution in both mono-static and bi-static sensing modes. The mean value characterizes the deterministic component, while the variance accounts for the stochastic component.
Different from the human body and UAV, vehicle RCS exhibit spiking characteristics that render them unsuitable for modeling with omnidirectional patterns. The RCS of humans and UAVs can be modeled using a log-normal distribution, in which the mean $\mu$ and variance $\sigma$ are determined from empirical measurements for random generation. displays four distinct peaks in its angular distribution, indicating the need for further exploration of its modeling and simulation methodologies.
$RCS_{Bi}=RCS_{Mono}\cos{\frac{\beta}{2}}$ (17)
This equation is valid under the condition that the target is sufficiently smooth and free from shadowing and backscatter effects. necessitates validation through measurement results and can be modeled analogous to an antenna radiation pattern.
3 Evaluation Metrics for ISAC Testing
3.1 Accuracy metric
Table 2: 3GPP ISAC Standardization Focused Scenarios
         Sensing usage          Principle          Accuracy          Resolution
         Ranging          $R=\frac{1}{2}ct_{R}$          $dR=\frac{1}{2}cdt_{R}+\frac{R}{C}dc$          $\Delta R=\frac{1}{2}ct_{\tau}$
         Speed detection          $f_{d}=\frac{2v}{\lambda}$          $dv=\frac{\lambda}{2}df_{d}$          $\Delta v=\frac{\lambda}{2}\Delta f_{d}$
         Angulation          $d\theta=\arcsin(\frac{\phi\lambda}{2\pi d})$          $d\theta=\frac{\lambda d\phi}{2\pi d\cos{\theta}}$          $\Delta\theta>\frac{\lambda}{Nd\cos{\theta}}$
The fundamental detection equation in radar is presented in Table 2. Here,$c$ is the speed of light with a value of $3\times 10^{8}m/s$ and $t_{R}$ represents the time difference between the transmitted pulse and the received echo pulse from the sensed target. The ranging accuracy can be qualitatively analyzed using the total differential expression of equation 19. The range resolution $\Delta R$ is defined as the minimum distinguishable distance, which is related to the pulse width $t_{\tau}$ of the sensing signal. The maximum unambiguous range is determined by the pulse repetition interval $T_{R}$ and is expressed as follows:
$R_{Max}=\frac{1}{2}cT_{R}$ (18)
$R=\frac{1}{2}c(mT_{R}+t_{r})$ (19)
The basic principle of angle measurement in the radar field is as follows.
$\phi=\frac{2\pi}{\lambda}d\sin{\theta}$ (20)
where $\phi$ represents the phase difference between different antenna elements that are separated by a distance $d$ at the receiver and $\theta$ is the angle of the sensing target. The angle measurement accuracy can be expressed as
$d\phi=\frac{2\pi}{\lambda}d\cos{\theta}d\theta$ (21)
$\phi=\frac{2\pi}{\lambda}d\sin{\theta}<\pi$ (22)
from equation 22,$\theta_{max}=\arcsin{(\frac{\lambda}{2\pi})}$. The angle resolution $\Delta\theta$ is defined as the minimum distinguishable angle, which is related to the main lobe width of the antenna array $\frac{2\pi}{N}$. according to LABEL:eq_ADresolution,$\Delta\theta>\frac{\lambda}{Nd\cos{\theta}}$ where $N$ is the number of antenna arrays. the basic principle of velocity measurement is based on the Doppler frequency. from the above equation,we obtain $v=\frac{\lambda f_{d}}{2}$.analogous to range resolution and angular resolution. Here, the Doppler resolution $\Delta f_{d}=\frac{1}{MT_{R}}$, in which $M$ is the number of transmitted pulses, and $T_{R}$ is the pulse repetition interval.
3.2 Detection metric
The probability of false alarm arises when the noise at the receiver surpasses the decision threshold, resulting in erroneous detection. This probability is associated with the root mean square (RMS) value of the noise and the decision threshold, and it is defined as follows
[…]
where $V_{T}$ is the decision threshold, $\omega(x)$ denotes the probability density function (PDF) of detection. When the noise is complex and both its real and imaginary components follow a Gaussian distribution, the magnitude of the noise follows a Rayleigh distribution,$\omega(x)$ can be expressed as
$\omega(x|\text{no target})=\frac{x}{\sigma^{2}}exp(-\frac{x^{2}}{2\sigma^{2}}),x\geq 0$ (24)
substituting Equation 24 into Equation 23 results in
$P_{fa}=exp(-\frac{V_{T}^{2}}{2\sigma^{2}})$ (25)
further, the relationship between $P_{fa}$ and $V_{T}$ is obtained as $V_{T}=\sqrt{2\sigma\ln{P_{fa}}}$ The probability of detection, $P_{D}$, represents the likelihood of correctly detecting a target when one is present and is defined as follows.
[…]
4 Simulation Results
In the simulation experiments, the primary focus is on the reliability of the ISAC channel cascading method and the accuracy of the RCS simulation results. To validate the channel cascading method, specific scenario parameters utilized in the dual-station sensing mode for a UAV scenario. The statistical characteristics of the generated channels—specifically, delay spread, horizontal angle of arrival (AoA) spread, horizontal angle of departure (AoD) spread, vertical AoA spread, and vertical AoD spread—are validated against the 5G standard channel model, the full convolution method, and the proposed parameter convolution method. The number of simulation runs is established at 1000 to collect statistical data on the spread values of the generated channel. To validate the RCS simulation results, the mean and variance of the RCS, based on the log-normal distribution, are derived by fitting the measured RCS results reported in document [15]. The RCS is simulated using the fitted distribution, and the simulated RCS results are compared with the measured data to assess the error.
Figure 3: Delay spread of ISAC concatenated Channel
As illustrated in Fig. 3, the delay spread of the cascaded channels consistently exceeds that of the standard channel. This phenomenon is attributed to the accumulation of the two delay components during the cascading process, which results in a broader overall delay distribution function and an increase in the spread value statistics. The spread value obtained through the parameter convolution method is smaller than that achieved with the full convolution method. This reduction is due to the parameter convolution approach, which decreases complexity by pruning clusters, thus eliminating low-power, high-delay clusters and subsequently reducing the spread value.
ASA of ISAC Channel
ASD of ISAC Channel
ESA of ISAC Channel
ESD of ISAC Channel
Figure 4: Angle spread of ISAC Channel
RCS Human 5GHz
RCS Human 14GHz
RCS Human 28GHz
RCS UAV 5GHz
RCS UAV 14GHz
RCS UAV 28GHz
Figure 5: RCS of UAV and Human in 5GHz, 14GHz, and 28GHz respectively
Fig. 4 demonstrates that the statistical values of the angle spread for the three channels closely align, particularly in the dimension of horizontal angle spread. This phenomenon arises from the cascading scheme, which generates statistical parameters for each segment. The mean and variance distributions remain consistent across simulations of the same scenario, resulting in the convergence of spread values through multiple statistical averages. A deviation is observed in 4, where the spread value of the cascaded scheme exceeds that of the standard channel. This phenomenon arises from the cluster pruning process retaining high-power clusters that are closely aligned with the angle of arrival. Consequently, the duplication of high-power clusters following pruning and cascading results in an increase in the spread value. The spread value associated with the parameter convolution method is lower than that of the full convolution method, attributed to the reduction of clusters with larger angles following pruning. Figure 5 illustrates the RCS simulation results for two target types: human and UAV, at frequencies of 5 GHz, 14 GHz, and 28 GHz. The simulated results demonstrate a strong alignment with the measured results reported in reference [14], thereby validating the feasibility of the log-normal fitting scheme. Additionally, this finding confirms the differences in RCS distribution between ISAC channels and traditional radar systems. This discrepancy arises from the fact that traditional radar RCS perception typically occurs at greater distances, while ISAC services are concentrated on closer ranges. Consequently, this focus results in variations between the RCS distributions observed in radar applications and those measured in ISAC.
5 Conclusion
In this paper, we extend the 3D-GBSM 5G standard channel model by incorporating the latest advancements in 3GPP ISAC standardization, thereby enhancing the integrated sensing and communication channel simulation capabilities for 6G. The new model accommodates both Bi-Static and Mono-Static sensing modes, supporting four types of sensing targets: UAVs, vehicles, humans, and AGVs. This model aligns with various options outlined in the 3GPP standardization progress, including methods for modeling RCS size and scale, polarization techniques, and power normalization for target and background channels.
By employing a parameter cascading method, we preserved the statistical characteristics of the ISAC channel distribution while achieving a 93.75% reduction in average convolution complexity. The simulation database for target RCS was developed by fitting actual measurements corresponding to the target RCS. A comparative analysis demonstrated a high degree of consistency between the simulated RCS and the measured RCS.
This advancement accelerates the standardization of the 6G ISAC channel model and the evaluation of novel algorithms, as it enables the utilization of existing 5G standard parameter tables alongside the proposed cascading model and measured RCS data to simulate ISAC channels. The modified simulation framework is applicable to all future developments in ISAC channel model standardization, offering insights into the standardization simulation process and facilitating the progress of simulation evaluations.
[…]
Appendix A.